Please help solve this equation (5+sqrt-64)-(23i-32)
Answer by algebrahouse.com(1659)

It's not an equation, but it is an expression that can be simplified:
(5 + √-64) - (23i - 32)
= 5 + √-1 √64 - 23i - 32 {broke √-64 into √-1 and √64}
= 5 + 8i - 23i - 32 {square root of -1 is i and square root of 64 is 8}
= -27 - 15i {combined like terms}
www.algebrahouse.com